OpenStudy (anonymous):

I am not good with word problems. :/ I need some help with this word problem. The area of a triangle is 30inches^2. Its base is 12in and its height is X inches. Write and solve an equation to find the value of X.

OpenStudy (anonymous):

area = 1/2 * base * height 30 = 1/2 * 12 * height 30/6= height so height= 5 in ans